# A simple algorithm for output range analysis for deep neural networks

### v3

@@ Abstract
This paper presents a simple algorithm for the output range estimation problem in Deep Neural Networks (DNNs) by integrating a Simulated Annealing (SA) algorithm tailored to operate within constrained domains and ensure convergence towards global optima. The method effectively addresses the challenges posed by the lack of local geometric information and the high non-linearity inherent to DNNs, making it applicable to a wide variety of architectures, with a special focus on Residual Networks (ResNets) due to their practical importance. Unlike existing methods, our algorithm imposes minimal assumptions on the internal architecture of neural networks, thereby extending its usability to complex models. Theoretical analysis guarantees convergence, while extensive empirical evaluations—including optimization tests involving functions with multiple local minima—demonstrate the robustness of our algorithm in navigating non-convex response surfaces. The experimental results highlight the algorithm’s efficiency in accurately estimating DNN output ranges, even in scenarios characterized by high non-linearity and complex constraints. For reproducibility, Python codes and datasets used in the experiments are publicly available through our GitHub repository ***https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
@@ 1 Introduction
Unquestionably, in recent decades, Deep Neural Networks (DNNs) have been by far the most widely used tools to perform complex machine learning tasks. More recently, DNNs have been used in cyber-physical systems critical to public security and integrity; such as autonomous vehicle driving and air traffic systems. Therefore, it is of pressing interest to implement security verification systems for DNNs. One of the objectives in this line of interest is the verification of the maximum and minimum values assumed by a DNN, an objective commonly known as the range estimation problem, see Dutta et al., (2018); Wang et al., (2018). This interest in estimating the range assumed by a DNN responds to the objective of diagnosing and validating the previously executed training. However, the relationships established between the inputs and outputs of a DNN are highly non-linear and complex, making it difficult to understand with existing tools today. Due to this inability, DNNs are commonly referred to as black boxes. This nature of DNN makes the range estimation problem particularly challenging because there is no geometric information about the response surface generated by a DNN. For example, if local geometric information about the generated surface by a DNN was obtained, such as the gradient vector and the Hessian matrix at each point, the problem could be addressed with conventional nonlinear programming techniques. However, in a DNN it is only possible to obtain point information about the estimated response, without any local knowledge around that point. These facts, added to the high nonlinearity of a DNN, make the range estimation problem nontrivial. In particular, this problem is even more challenging when dealing with a DNN, in contrast to other types of neural network, since its multiple layers increase the complexity of the problem. In this paper, we present a simple algorithm based on global optimization techniques, initially motivated by the classical Simulated Annealing (SA) kirkpatrick1983optimization, to solve the range estimation problem for a wide spectrum of neural networks, and in particular DNNs. In contrast to classical SA, our algorithm considers the existence of a restricted domain for searching for optimal points, which eventually corresponds to the domain of the training data of the analyzed neural network. Our algorithm does not make use of any information about the internal architecture of the analyzed DNN, which makes it applicable to a much larger spectrum than other proposals available in the literature, see for example Dutta et al., (2018); Wang et al., (2018); katz2017reluplex. Our algorithm considers restricted search spaces, which underlie the nature of the output range estimation problem of neural networks. Furthermore, we present results, both theoretical and empirical, that guarantee the convergence of our algorithm towards the optimal points, which leads to a good estimation of the output range.
@@ 2 Related Work
Recently, various methodologies have been proposed for the analysis of the output range of DNNs. Dutta et al., (2018) propose a novel approach to estimate the output ranges of these models given specific input sets. Their novel algorithm utilizes local search techniques and linear programming to efficiently compute the maximum and minimum values in a DNN over the input set. The algorithm repeatedly applies local gradient descent to identify and eliminate local minima of the neural network’s output function. Once local optima are identified, the final global optimum is validated through a mixed integer programming model. The authors consider the algorithm’s computational efficiency while still maintaining the accuracy of the output range estimation. However, common issues like scalability and the potential complexity of a DNN are not fully addressed. Furthermore, the methodology makes specific assumptions about the neural network, which means that the methodology can only be applied to a rather restricted class of neural networks.
@@
On the other hand, autonomous vehicles, collision avoidance systems, and others are part of a real-world security-critical domain, making it important to check the security properties of DNNs. These security properties guarantee that a DNN should have to ensure it operates safely and reliably against potential threats. There are frameworks for verifying the security properties of DNN that use interval arithmetic Wang et al., (2018); katz2017reluplex; carlini2018provably. Wang et al., (2018) shows us a framework called ReluVal, which employs interval arithmetic to represent the ranges of inputs and outputs. For each input and operation in a DNN, the interval calculation is performed to explore the output range of the network. By analyzing these intervals, the framework can detect inputs that lead to incorrect classifications or outputs, thereby exposing potential weaknesses in the network. The experiments demonstrate the efficiency of ReluVal in verifying security properties, outperforming other frameworks. Nevertheless, it may still struggle with scalability when applied to extensive networks or complex architectures. Furthermore, the framework is focused on specific types of adversarial attacks, leaving uncertainty regarding its performance on different architectures.
@@
Other kinds of verification of DNNs are required: safety and robustness. Tran et al., (2020) presents the Neural Network Verification (NNV) software tool. The tool applies a collection of reachability algorithms and the use of a variety of set representations, such as polyhedra, star sets, zonotopes, and abstract-domain representations. Also, NNV can handle different types of neural networks and system models, allowing it to support both exact and over-approximate analysis. However, representations as polyhedra are limited by scalability and may not work effectively, as they involve costly operations, leading to a conservative and less practical reachability analysis in larger systems. Moreover, all experiments focus on specific exercises, which limits the exploration of new and complex DNN architectures. Then, Liu et al., (2021) provide a comprehensive overview of methods developed to formally verify the safety and robustness of DNNs. The authors discuss different algorithms that incorporate reachability analysis, optimization, and search techniques to verify the appropriate configuration of the DNNs for the input-output properties across their entire input space. These algorithms are classified into different categories, such as layer-by-layer, reachability analysis, optimization-based, and combined search and verification methods. However, after exhaustive verification of these algorithms, the work concludes with some algorithms that present, like the other works, problems with the network scale. Finally, Huang et al., (2020) analyze the output range of a neural network using a convex polygonal relaxation (over-approximation) of the activation functions to manage nonlinearity, allowing the problem to be formulated as a mixed-integer linear program (MILP). However, a key limitation is the addition of more integer variables in the MILP, which increases computational complexity, particularly in DNNs or those with more complex architectures.
@@ 3 Residual Neural Networks
As mentioned above, since we make no assumptions about the internal architecture of the analyzed neural network, our algorithm works for a wide spectrum of neural networks. However, without loss of generality, we focus on Residual Neural Networks given their relevance in applications and for illustrative purposes of the use of our algorithm. The Residual Networks (or ResNets) is a kind of DNN that was introduced in He et al., (2016), for which in this chapter we show a mathematical description. Let $x\in\mathbb{R}^{d}$ denote the inputs and $y\in\mathbb{R}$ be the output in a supervised learning problem. In this learning context, the objective is the approximation of a function $f:x\in\mathbb{R}^{d}\mapsto y\in\mathbb{R}$ using a ResNet, which we denote as $\mathcal{F}$. The basic unitary components of a ResNet are called residual blocks, which are organized in a consecutive number of layers that are linked through nonlinear functions. To obtain a mathematical description of the operations carried out within a ResNet, we define some notations. We consider a network with a source layer, a output layer and $L$ hidden layers, where the $l$-th layer contains $H_{l}$ neurons ($l$-th layer width), for $l=0,\dots,L+1$. Note that $H_{0}=d$ and $H_{L+1}=1$, corresponding to the feature inputs and target outputs respectively. We denote the output vector of the $l$-th layer by $x^{(l)}\in\mathbb{R}^{H_{l}}$ which corresponds to the input of the next layer. We set $x^{(0)}=x\in\mathbb{R}^{d}$ and $x^{(L+1)}=y\in\mathbb{R}$. For $l=1,\dots,L$, the $i$-th neuron performs an affine transformation on that layers input $x^{(l-1)}$ followed by a non-linear transformation
$x^{(l)}_{i}=\sigma\Bigg{(}\sum\limits_{j=1}^{H_{l-1}}W_{ij}x^{(l-1)}_{j}+b_{i}^{(l)}\Bigg{)}+x^{(l-1)}_{i},\quad 1\leq i\leq H_{l},$ (1)
@@
$x^{(l)}=\sigma\Big{(}\mathcal{A}^{(l)}\big{(}x^{(l-1)}\big{)}\Big{)}+x^{(l-1)},\qquad\mathcal{A}^{(l)}\big{(}x^{(l-1)}\big{)}=W^{(l)}x^{(l-1)}+b^{(l)},$ (2)
@@
Although the activation function can, in principle, be chosen arbitrarily, there are three functions that have particularly proven to be useful in various applications; ReLU, Sigmoid, and Tanh, see LeCun et al., (2015); Goodfellow et al., (2016). The parameters of the network is all the weights and biases $\{W^{(l)},b^{(l)}\}_{l=1}^{L+1}\in\mathbb{R}^{N_{p}}$, where $N_{p}=\sum\limits_{l=1}^{L+1}(H_{l-1}+1)H_{l}$ is the total number of parameters.
@@ 4 Output Range Analysis Problem
In the rest of this paper, we consider a ResNet $\mathcal{F}$ with inputs $x\in\mathbb{R}^{d}$ and output $y\in\mathbb{R}$ as defined (3). We assume that the network parameters of the neural network were optimally estimated following standard training procedures such as those shown in Bishop, (2006); Murphy, (2012).
@@
    OUTPUT: An interval $[\mathcal{F}_{min},\mathcal{F}_{max}]$ such that $\mathcal{F}(x)\in[\mathcal{F}_{min},\mathcal{F}_{max}]$, i.e., $[\mathcal{F}_{min},\mathcal{F}_{max}]$ contains the range of $\mathcal{F}$ over inputs $x\in E$. Moreover, the optimal points $\{x_{min},x_{max}\}\in E$ such that $\mathcal{F}(x_{min})=\mathcal{F}_{min}$ and $\mathcal{F}(x_{max})=\mathcal{F}_{max}$.
@@
Due to the applications, we focus just on hypercubes of $\mathbb{R}^{d}$ as feasible set, that is; sets of the type $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $[l_{j},u_{j}]$ are intervals in $\mathbb{R}$. Note that $E$ is the restricted search domain in which we want to diagnose and validate the analyzed neural network.

### v4

@@ Abstract
This paper presents a simple algorithm for the output range estimation problem in Deep Neural Networks (DNNs) by integrating a Simulated Annealing (SA) algorithm tailored to operate within constrained domains and ensure convergence towards global optima. The method effectively addresses the challenges posed by the lack of local geometric information and the high non-linearity inherent to DNNs, making it applicable to a wide variety of architectures, with a special focus on Residual Networks (ResNets) due to their practical importance. Unlike existing methods, our algorithm imposes minimal assumptions on the internal architecture of neural networks, thereby extending its usability to complex models. Theoretical analysis guarantees convergence, while extensive empirical evaluations—including optimization tests involving functions with multiple local minima—demonstrate the robustness of our algorithm in navigating non-convex response surfaces. The experimental results highlight the algorithm’s efficiency in accurately estimating DNN output ranges, even in scenarios characterized by high non-linearity and complex constraints.
@@ 1 Introduction
Unquestionably, in recent decades, Deep Neural Networks (DNNs) have been by far the most widely used tools to perform complex machine learning tasks. More recently, DNNs have been used in cyber-physical systems critical to public security and integrity; such as autonomous vehicle driving and air traffic systems. Therefore, it is of pressing interest to implement security verification systems for DNNs. One of the objectives in this line of interest is the verification of the maximum and minimum values assumed by a DNN, an objective commonly known as the range estimation problem, see Dutta et al. (2018); Wang et al. (2018). This interest in estimating the range assumed by a DNN responds to the objective of diagnosing and validating the previously executed training. However, the relationships established between the inputs and outputs of a DNN are highly non-linear and complex, making it difficult to understand with existing tools today. Due to this inability, DNNs are commonly referred to as black boxes. This nature of DNN makes the range estimation problem particularly challenging because there is no geometric information about the response surface generated by a DNN. For example, if local geometric information about the generated surface by a DNN was obtained, such as the gradient vector and the Hessian matrix at each point, the problem could be addressed with conventional nonlinear programming techniques. However, in a DNN it is only possible to obtain point information about the estimated response, without any local knowledge around that point. These facts, added to the high nonlinearity of a DNN, make the range estimation problem nontrivial. In particular, this problem is even more challenging when dealing with a DNN, in contrast to other types of neural network, since its multiple layers increase the complexity of the problem. In this paper, we present a simple algorithm based on global optimization techniques, initially motivated by the classical Simulated Annealing (SA) Kirkpatrick et al. (1983), to solve the range estimation problem for a wide spectrum of neural networks, and in particular DNNs. In contrast to classical SA, our algorithm considers the existence of a restricted domain for searching for optimal points, which eventually corresponds to the domain of the training data of the analyzed neural network. Our algorithm does not make use of any information about the internal architecture of the analyzed DNN, which makes it applicable to a much larger spectrum than other proposals available in the literature, see for example Dutta et al. (2018); Wang et al. (2018); Katz et al. (2017). Our algorithm considers restricted search spaces, which underlie the nature of the output range estimation problem of neural networks. Furthermore, we present results, both theoretical and empirical, that guarantee the convergence of our algorithm towards the optimal points, which leads to a good estimation of the output range.
In line with the objective and motivation stated—estimating, in a reliable way, the extreme values that the network can produce under bounded inputs while treating it as a black box—we adopt a global search strategy on a finite domain that combines four simple and complementary ideas. First, symmetric random proposals are generated around the current point to explore without directional bias. Second, a cyclic reflection is applied at the domain boundaries: any proposal that leaves the range “bounces” back in by symmetry, preventing artificial accumulation near the boundary. Third, a probabilistic acceptance criterion balances exploration and exploitation and, in a controlled way, allows accepting moves that do not immediately improve in order to escape local optima. Fourth, the intensity of exploration is reduced gradually so that the search progressively concentrates on the most promising regions. These elements are articulated into a single update mechanism: at each iteration a candidate is proposed (with reflection if needed), the network output is evaluated at that point, and acceptance or rejection is decided; if accepted, the state is updated, and in all cases the best value found and its location are recorded. The formal details and assumptions are presented in the methodology Section 5. The rationale for this dynamics is that it combines broad, unbiased coverage with progressive focusing. Symmetric proposals together with cyclic reflection support a balanced exploration of the admissible space, preventing the trajectory from sticking to the boundaries and encouraging visits to distinct regions of the domain. The probabilistic acceptance criterion acts as a safeguard against entrapment: early on it facilitates barrier crossing and exits from local optima; later, as the exploration intensity is gradually reduced, it acts as an adaptive “zoom” that concentrates sampling where the evidence is stronger. Moreover, by always retaining the best value observed, the procedure does not lose progress: even while exploring alternatives, the best estimate of the extreme improves monotonically. In a black-box setting with bounded inputs, this sequence makes visits to neighborhoods of the true extrema very likely and, under mild continuity, yields estimates close to the extreme values of interest.
Our methodology for output range estimation in deep neural networks shares similarities with global optimization methods used in previous studies. For example, Gromov et al. Gromov et al. (2024) and Kovaleva & Smirnov Kovaleva and Smirnov (2024) explore optimization techniques for finding global extrema in complex systems, including neural networks. These methods, such as Simulated Annealing (SA) and memetic algorithms, are effective for general optimization problems. However, while these studies focus on optimizing various types of systems, our work introduces a novel approach by applying SA with reflective boundary conditions specifically designed for deep neural networks with constrained input spaces. This reflects a need in the literature for methods that preserve boundaries while optimizing over complex, high-dimensional spaces. Furthermore, previous works like the one by van Laarhoven & Aarts van Laarhoven and Aarts (1987) and Nourani & Andresen Nourani and Andresen (1998) provide strong theoretical foundations for SA in global optimization. Our approach complements these methods by maintaining asymptotic convergence guarantees while incorporating constraints imposed by the architecture of deep neural networks. Additionally, recent studies such as the one by Gromov et al. Gromov et al. (2024) on Bayesian optimization and the study of evolutionary algorithms for global optimization by Kovaleva & Smirnov Kovaleva and Smirnov (2024) further emphasize the importance of global extrema search in machine learning contexts. Our method extends these approaches by directly addressing the challenges posed by the constrained domains in neural network outputs, offering a practical solution for deep learning models. Finally, optimization methods and extremal problems in neural network training, such as those explored by Gromov et al. Gromov et al. (2024), are directly addressed by our technique. While previous studies have applied evolutionary strategies and optimization algorithms to explore global minima, our approach with reflective boundaries adds a layer of robustness by ensuring that the optimization process remains within feasible input domains, thus overcoming the limitations of traditional methods in constrained environments.
@@ 2 Related Work
Recently, various methodologies have been proposed for the analysis of the output range of DNNs. Dutta et al. (2018) propose a novel approach to estimate the output ranges of these models given specific input sets. Their novel algorithm utilizes local search techniques and linear programming to efficiently compute the maximum and minimum values in a DNN over the input set. The algorithm repeatedly applies local gradient descent to identify and eliminate local minima of the neural network’s output function. Once local optima are identified, the final global optimum is validated through a mixed integer programming model. The authors consider the algorithm’s computational efficiency while still maintaining the accuracy of the output range estimation. However, common issues like scalability and the potential complexity of a DNN are not fully addressed. Furthermore, the methodology makes specific assumptions about the neural network, which means that the methodology can only be applied to a rather restricted class of neural networks.
@@
On the other hand, autonomous vehicles, collision avoidance systems, and others are part of a real-world security-critical domain, making it important to check the security properties of DNNs. These security properties guarantee that a DNN should have to ensure it operates safely and reliably against potential threats. There are frameworks for verifying the security properties of DNN that use interval arithmetic Wang et al. (2018); Katz et al. (2017); Carlini et al. (2018). Wang et al. (2018) shows us a framework called ReluVal, which employs interval arithmetic to represent the ranges of inputs and outputs. For each input and operation in a DNN, the interval calculation is performed to explore the output range of the network. By analyzing these intervals, the framework can detect inputs that lead to incorrect classifications or outputs, thereby exposing potential weaknesses in the network. The experiments demonstrate the efficiency of ReluVal in verifying security properties, outperforming other frameworks. Nevertheless, it may still struggle with scalability when applied to extensive networks or complex architectures. Furthermore, the framework is focused on specific types of adversarial attacks, leaving uncertainty regarding its performance on different architectures.
@@
Other kinds of verification of DNNs are required: safety and robustness. Tran et al. (2020) presents the Neural Network Verification (NNV) software tool. The tool applies a collection of reachability algorithms and the use of a variety of set representations, such as polyhedra, star sets, zonotopes, and abstract-domain representations. Also, NNV can handle different types of neural networks and system models, allowing it to support both exact and over-approximate analysis. However, representations as polyhedra are limited by scalability and may not work effectively, as they involve costly operations, leading to a conservative and less practical reachability analysis in larger systems. Moreover, all experiments focus on specific exercises, which limits the exploration of new and complex DNN architectures. Then, Liu et al. (2021) provide a comprehensive overview of methods developed to formally verify the safety and robustness of DNNs. The authors discuss different algorithms that incorporate reachability analysis, optimization, and search techniques to verify the appropriate configuration of the DNNs for the input-output properties across their entire input space. These algorithms are classified into different categories, such as layer-by-layer, reachability analysis, optimization-based, and combined search and verification methods. However, after exhaustive verification of these algorithms, the work concludes with some algorithms that present, like the other works, problems with the network scale. Finally, Huang et al. (2020) analyze the output range of a neural network using a convex polygonal relaxation (over-approximation) of the activation functions to manage nonlinearity, allowing the problem to be formulated as a mixed-integer linear program (MILP). However, a key limitation is the addition of more integer variables in the MILP, which increases computational complexity, particularly in DNNs or those with more complex architectures.
@@ 3 Residual Neural Networks
As mentioned above, since we make no assumptions about the internal architecture of the analyzed neural network, our algorithm works for a wide spectrum of neural networks. However, without loss of generality, we focus on Residual Neural Networks given their relevance in applications and for illustrative purposes of the use of our algorithm. The Residual Networks (or ResNets) is a kind of DNN that was introduced in He et al. (2016), for which in this chapter we show a mathematical description. Let $x\in\mathbb{R}^{d}$ denote the inputs and $y\in\mathbb{R}$ be the output in a supervised learning problem. In this learning context, the objective is the approximation of a function $f:x\in\mathbb{R}^{d}\mapsto y\in\mathbb{R}$ using a ResNet, which we denote as $\mathcal{F}$. The basic unitary components of a ResNet are called residual blocks, which are organized in a consecutive number of layers that are linked through nonlinear functions. To obtain a mathematical description of the operations carried out within a ResNet, we define some notations. We consider a network with a source layer, a output layer and $L$ hidden layers, where the $l$-th layer contains $H_{l}$ neurons ($l$-th layer width), for $l=0,\dots,L+1$. Note that $H_{0}=d$ and $H_{L+1}=1$, corresponding to the feature inputs and target outputs respectively. We denote the output vector of the $l$-th layer by $x^{(l)}\in\mathbb{R}^{H_{l}}$ which corresponds to the input of the next layer. We set $x^{(0)}=x\in\mathbb{R}^{d}$ and $x^{(L+1)}=y\in\mathbb{R}$. For $l=1,\dots,L$, the $i$-th neuron performs an affine transformation on that layers input $x^{(l-1)}$ followed by a non-linear transformation
$x^{(l)}_{i}=\sigma\Bigg(\sum_{j=1}^{H_{l-1}}W_{ij}x^{(l-1)}_{j}+b_{i}^{(l)}\Bigg)+x^{(l-1)}_{i},\quad 1\leq i\leq H_{l}.$ (1)
@@
$\begin{split}x^{(l)}&=\sigma\Big(\mathcal{A}^{(l)}\big(x^{(l-1)}\big)\Big)+x^{(l-1)},\\[6.0pt] \mathcal{A}^{(l)}\big(x^{(l-1)}\big)&=W^{(l)}x^{(l-1)}+b^{(l)}.\end{split}$ (2)
@@
Although the activation function can, in principle, be chosen arbitrarily, there are three functions that have particularly proven to be useful in various applications; ReLU, Sigmoid, and Tanh, see LeCun et al. (2015); Goodfellow et al. (2016). The parameters of the network is all the weights and biases $\{W^{(l)},b^{(l)}\}_{l=1}^{L+1}\in\mathbb{R}^{N_{p}}$, where $N_{p}=\sum\limits_{l=1}^{L+1}(H_{l-1}+1)H_{l}$ is the total number of parameters.
Residual blocks with general activations.
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a Lipschitz activation (e.g., ReLU, $\tanh$, GELU). To make the residual path explicit and avoid ambiguity between (2) and (3), we write
$x^{(0)}$ $=x,$
$x^{(l+1)}$ $=\sigma\!\big(A^{(l)}(x^{(l)})\big)+x^{(l)},\qquad l=0,\ldots,L-1.$
and the network output
$\mathcal{F}(x)=\mathcal{A}^{(L+1)}\!\big(x^{(L)}\big).$
These expressions preserve the skip connection for any choice of $\sigma$. Note that our algorithm treats the network as a black box; the presence of residual connections is not required by the method or by the analysis. In this case, we display the residual formulation to adopt notation that is standard in modern architectures and to align with the models used in our experiments, ResNet-based surrogates of Ackley and Drop–Wave. The simulated-annealing procedure and the theoretical development, however, treat $\mathcal{F}$ as a black box; the same analysis applies verbatim to generic feedforward DNNs with affine layers and a Lipschitz activation.
@@ 4 Output Range Analysis Problem
In the rest of this paper, we consider a ResNet $\mathcal{F}$ with inputs $x\in\mathbb{R}^{d}$ and output $y\in\mathbb{R}$ as defined (3). We assume that the network parameters of the neural network were optimally estimated following standard training procedures such as those shown in Bishop (2006); Murphy (2012).
@@
    OUTPUT: An interval $\mathcal{F}_{min},\mathcal{F}_{max}]$ such that $\mathcal{F}(x)\in[\mathcal{F}_{min},\mathcal{F}_{max}]$, i.e., $\mathcal{F}_{min},\mathcal{F}_{max}]$ contains the range of $\mathcal{F}$ over inputs $x\in E$. Moreover, the optimal points $\{x_{min},x_{max}\}\in E$ such that $\mathcal{F}(x_{min})=\mathcal{F}_{min}$ and $\mathcal{F}(x_{max})=\mathcal{F}_{max}$.
@@
Due to the applications, we focus just on hypercubes of $\mathbb{R}^{d}$ as feasible set, that is; sets of the type $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $l_{j},u_{j}]$ are intervals in $\mathbb{R}$. Note that $E$ is the restricted search domain in which we want to diagnose and validate the analyzed neural network.